OpenStudy (anonymous):

How do you integrate sin (x^2)?

OpenStudy (unklerhaukus):

\[\int\sin(x^2)\,\mathrm dx\]

OpenStudy (unklerhaukus):

let \(u=x^2\) \(u^{1/2}=x\) \(\tfrac12u^{-1/2}\,\mathrm du=\mathrm dx\)

OpenStudy (unklerhaukus):

hmm, it doesn't really seam to work out nice, are you sure your question wasn't ∫sin^2(x) dx ?

OpenStudy (anonymous):

Yes, I'm sure- that's how the question was framed. Maybe it was a mistake from the examiner's end?

OpenStudy (unklerhaukus):

\[=\frac12\int u^{-1/2}\sin(u)\,\mathrm du\]

OpenStudy (unklerhaukus):

@sauravshakya